Created by Main.MirkoPojer on 2007-06-29 - 20:33
D2-Q4 study in the afternoon
The following cycle has been executed on D2-Q4, to study stability at minimum liquid helium level in DFBMA:
- with nominal liquil helium level inside the DFBMA, RD2.L8 and RQ4.L8B1/B2 to I_min_op
- lower liquid helium level to 31%
- ramp of RD2.L8 to 6kA with 18 A/s ramp rate; once at nominal, ramp of RQ4.L8B1/B2 to 3610 A with 10A/s ramp rate
- plateau of 30 min, then ramp RQ4.L8B1/B2 to I_min_op (RD2.L8 always at nominal)
- wait 2 min and ramp RQ4.L8B2 to 3610A at 10 A/s and RQ4.L8B1 to 1900A at 10A/s
- plateau of 20 min
- power permit removed on all PCs and SPA initiated withou quench.
Test result: successful!
Edit, attach file or add comment
Created by Main.AntonioVergara on 2007-06-29 - 17:16

PlanningW27.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x29x16x22/PlanningW27.xls
Comments:
Edit, attach file or add comment
Created by Main.MatteoSolfaroli on 2007-06-29 - 16:33
Activities in sector 45 from Monday 2nd July
- CRYO
- Purging on Friday 29th evening and all the week end long;
- Refilling on Monday 2nd July;
- ElQA-TP4B and Monitoring installation
- On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB) maybe from 8am to 1pm. RESTRICTED ACCESS.
- CRYO-instrumentation
- Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
- Commissioning;
- Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
- TS/EL intervention
- Changing of the connectors of a BPM (LSS4R);
- Connection check of the instrumentation cables of BCT;
- TIM camera
- General check on Monday 2nd July;
- Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
- Safety Inspection
- AT/ACR Patrouille
- A Patrouille will be carried out on Wednesday 4th July just before the cooldown.
- AT/VAC
- Short access to LSS4R and LSS5L on Monday 2nd July.
- AB/BI
- LSS4R short intervention on the BI (capeau).
Comments:
Edit, attach file or add comment